# RaCFormer: Towards High-Quality 3D Object Detection via Query-based Radar-Camera Fusion

### v1

@@ Abstract
We propose Radar-Camera fusion transformer (RaCFormer) to boost the accuracy of 3D object detection by the following insight. The Radar-Camera fusion in outdoor 3D scene perception is capped by the image-to-BEV transformation–if the depth of pixels is not accurately estimated, the naive combination of BEV features actually integrates unaligned visual content. To avoid this problem, we propose a query-based framework that enables adaptively sample instance-relevant features from both the BEV and the original image view. Furthermore, we enhance system performance by two key designs: optimizing query initialization and strengthening the representational capacity of BEV. For the former, we introduce an adaptive circular distribution in polar coordinates to refine the initialization of object queries, allowing for a distance-based adjustment of query density. For the latter, we initially incorporate a radar-guided depth head to refine the transformation from image view to BEV. Subsequently, we focus on leveraging the Doppler effect of radar and introduce an implicit dynamic catcher to capture the temporal elements within the BEV. Extensive experiments on nuScenes and View-of-Delft (VoD) datasets validate the merits of our design. Remarkably, our method achieves superior results of 64.9% mAP and 70.2% NDS on nuScenes, even outperforming several LiDAR-based detectors. RaCFormer also secures the 1st ranking on the VoD dataset. The code will be released.
@@ 1 Introduction
Figure 1: Motivation of RaCFormer. (a) Previous approaches typically fuse entire BEV features, from image-view transformation and radar point cloud encoding, by concatenation. (b) RaCFormer, on the other hand, uses a query-based fusion framework by simultaneously sampling radar-enhanced image-view features, camera-transformed BEV features, and radar-encoded BEV features.
@@
Despite impressive advancement, it is still a non-trivial problem in approaching the detection accuracy of LiDAR-based methods by the combination of camera and radar data. Current top-performing radar-camera fusion approaches [53, 17, 28, 52] typically adopt the bird-eye-view-based (BEV) fusion framework [33] which unifies the representation in BEV space to facilitate fusion. In this paradigm, image and radar features are independently extracted, unified into a BEV representation, and subsequently fused by concatenation, as shown in Fig. 1 (a). However, it is difficult to merge the disparity between these two modalities by simple concatenation. The radar’s limited number of antennas results in a low-resolution beam, thus producing extremely sparse radar BEV features. On the other hand, camera BEV features are generated from dense image features, but have the issue of feature distortion due to the inaccurate depth estimation in view transformation [42]. In contrast, the original perspective image features offer a semantically rich representation free from distortions, which has the potential to aid camera-radar fusion. This underscores the necessity of heterogeneous feature fusion across perspectives of the front view and BEV. Inspired by this finding, this work aims to explore an effective cross-perspective fusion framework that can accommodate resolution and semantic discrepancies between the two modalities.
@@
Formally, we propose RaCFormer, a query-based radar-camera fusion framework that improves camera-radar fusion by sampling object-relevant features from both perspective and bird’s-eye views, as illustrated in the dual-view fusion paradigm in Fig. 1 (b). Our framework has three main designs: linearly increasing circular query initialization, radar-aware depth prediction, and temporal radar BEV encoding. The first optimizes the initialization distribution of queries, while the latter two refine and bolster the BEV representation. Specifically, we propose a circular query initialization strategy that places query points along concentric circles to align the projection principle of sensors. Additionally, we ensure a linear increase in the number of queries from inner to outer circles, thereby mitigating the issue of queries being much sparser at distant ranges compared to nearby areas. Furthermore, radar points lack accurate height information relative to the ground. Therefore, we assign a default height and back-project them onto the image features to enhance depth prediction for view transformation, refining the camera BEV features. Furthermore, we utilize the radar’s Doppler effect to track object velocities by employing an implicit dynamic catcher with convolutional gated recurrent units, effectively capturing temporal elements on multi-frame radar BEV features.
@@
  • We perform experiments on the nuScenes and VoD datasets. Our method achieves the state-of-the-art performance on the nuScenes test set. On the VoD dataset, the performance of RaCFormer also ranks first.
@@ 3.1 Overall Framework
RaCFormer, as shown in Fig. 2, offers a query-based 3D object detection framework that integrates radar and camera inputs. The core modules of the framework include an image encoder, a pillar encoder, a radar-aware depth head, an LSS view transformation module, an implicit dynamic catcher, and a transformer decoder. The image encoder extracts features from camera frames, while the pillar encoder processes radar points and flattens the features to BEV. Subsequently, the radar BEV features are refined by an implicit dynamic catcher to capture moving elements. Radar points are also back-projected into the image plane and combined with visual features in the radar-aware depth head to form a depth probability distribution $D^{\prime}$. The enhanced depth distribution is merged with image features in the LSS module to generate camera BEV features. Queries serve as a medium for cross-perspective and cross-modality feature fusion, initialized in an adjustable circular distribution and refined by the transformer decoder. The transformer decoder, comprising $L$ layers, includes a scale-adaptive self-attention module [30] for dynamically adjusting receptive fields, two ray-sampling modules for extracting BEV and image-view features, and an adaptive mixer [8] for feature aggregation. Finally, the classification and regression heads interpret the refined queries for accurate object detection.
@@ 3.2 Camera-transformed BEV Feature Generation with Radar-aware Depth Prediction
Radar-aware Depth Prediction: We propose to enhance the image features using radar data for better depth estimation. However, the limitation of vehicle-mounted radars is that they provide distance and velocity data along the beam’s direction but do not directly offer accurate height information. As depicted in Fig. 3(a), many radar points are projected onto the image with their vertical coordinates falling outside the objects’ 2D bounding boxes, due to inaccuracies in the z-coordinates of the raw radar points. Therefore, we design a pre-processing step before the depth head, shown in Fig. 3(b). To maximize the number of radar points falling onto the image’s field of view, we initially set $z_{r}=1$ for all points, denoted as $(x_{r},y_{r},z_{r})$, and then project them onto the image plane using the camera’s intrinsic parameters. The specific transformation formula is as follows:
@@ 4.2 Implementation Details
Table 2: Comparison on the nuScenes test set. The VoVNet-99 (V2-99) [20] is pre-trained from DD3D [41] with extra data. ‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and radar, respectively. “(+$t$)” indicates using future and historical frames, each by $t$ frames.
@@ 4.4 Ablation Studies
Feature Decoding and Fusion: Tab. 4 illustrates the impact of feature decoding methods and perspective selection. Experiment (A) employs the standard BEVFusion paradigm with a center-head decoder. Switching to a transformer decoder with queries in Experiment (B), focusing solely on the sampling of BEV features, yields a 4.6% enhancement in mAP and a 4.9% improvement in NDS, along with a significant reduction in mAVE but an increase in mATE. Experiment (C) further samples image-view features, improving 3.1% at mAP and 4.1% at NDS, respectively.
Radar-aware Depth Prediction: In Tab. 5, we examine the impact of embedding radar points’ depth and RCS value. The model (ID=B) that incorporates radar depth embedding achieves an improvement of 0.7% in mAP and 1.2% in NDS compared to the model (ID=A) without any embedding. Similarly, the embedding of RCS (ID=C) enhances 0.5% mAP and 0.8% NDS. When both embeddings are used, the model performance (ID=D) further improves, with overall gains of 1.1% in mAP and 2.1% in NDS.
@@
A C B 35.9 42.7 0.625 0.289 0.613
B Q B 40.5 47.6 0.655 0.281 0.365
C Q B+I 43.6 51.7 0.577 0.274 0.341
Table 4: Ablation study of the feature decoding. ‘F’ denotes fusion methods, with ‘C’ and ‘Q’ indicating concatenation and query-based sampling, respectively. ‘V’ represents the view, with ‘B’ and ‘I’ corresponding to the BEV and image view, respectively.
Implicit Dynamic Catching: In Tab. 6, we evaluate the performance of our implicit dynamic catcher (IDC) by evaluating mAP and mAVE for moving objects in the nuScenes validation set, divided by objects’ velocity: static (0 m/s), slow ($<$5 m/s), and fast ($>$5 m/s) objects. Adding radar data without IDC substantially increases the relative mAP of slow and fast objects in model (A) by 49.2% and 48.1%, while decreasing relative mAVE by 30.2% and 11.8%. With IDC, RaCFormer further improves relative mAP by 4.3% and 3.5% and relatively reduces mAVE by 2.9% and 2.0%. The integration of radar data and the IDC module reveals subtle temporal feature changes, leading to more pronounced optimization on slow-moving objects.
@@
A 43.6 51.7 0.577 0.274 0.341
B 44.3 52.9 0.560 0.268 0.319
C 44.1 52.5 0.564 0.269 0.332
D 44.7 53.8 0.558 0.270 0.313
Table 5: Ablation study about the radar-aware depth prediction. “DE” and “RE” denote the depth embedding and RCS embedding of radar points in the pre-processing step of depth head.
ID Input IDC 0 m/s (0, 5] m/s $>$5 m/s
@@
Table 9: Analysis of robustness using Car class AP. “All” denotes that the single modality is entirely off.
Inference Time: To enable real-time detection, we develop a lightweight version of our model: it uses 4 historical frames, 450 queries, and a BEV grid of $64\times 64$ with a resolution of 1.6 meters per voxel. These adjustments halve the computational load compared to our standard model. Despite these reductions, our model still delivers state-of-the-art mAP and NDS of 51.0% and 58.8%, respectively, outperforming HyDRa [47] by 1.6% in mAP and 2.8% in NDS. Operating on a single RTX 3090 GPU, it achieves a frame rate of 12 FPS, satisfying real-time requirements.

### v2

@@ Abstract
We propose Radar-Camera fusion transformer (RaCFormer) to boost the accuracy of 3D object detection by the following insight. The Radar-Camera fusion in outdoor 3D scene perception is capped by the image-to-BEV transformation–if the depth of pixels is not accurately estimated, the naive combination of BEV features actually integrates unaligned visual content. To avoid this problem, we propose a query-based framework that enables adaptive sampling of instance-relevant features from both the bird’s-eye view (BEV) and the original image view. Furthermore, we enhance system performance by two key designs: optimizing query initialization and strengthening the representational capacity of BEV. For the former, we introduce an adaptive circular distribution in polar coordinates to refine the initialization of object queries, allowing for a distance-based adjustment of query density. For the latter, we initially incorporate a radar-guided depth head to refine the transformation from image view to BEV. Subsequently, we focus on leveraging the Doppler effect of radar and introduce an implicit dynamic catcher to capture the temporal elements within the BEV. Extensive experiments on nuScenes and View-of-Delft (VoD) datasets validate the merits of our design. Remarkably, our method achieves superior results of 64.9% mAP and 70.2% NDS on nuScenes. RaCFormer also secures the state-of-the-art performance on the VoD dataset. Code is available at https://github.com/cxmomo/RaCFormer.
@@ 1 Introduction
Figure 1: Motivation of RaCFormer. (a) Previous methods typically fuse BEV features from image-view transformation and radar point cloud encoding, by concatenation or cross-attention. (b) Instead, RaCFormer uses a query-based fusion framework by simultaneously sampling radar-enhanced image-view features, camera-transformed BEV features, and radar-encoded BEV features.
@@
Despite impressive advancement, it is still a non-trivial problem in approaching the detection accuracy of LiDAR-based methods by the combination of camera and radar data. Current top-performing radar-camera fusion approaches [54, 18, 29, 53] typically adopt the BEV-based fusion framework [34] which unifies the representation in BEV space to facilitate fusion. In this paradigm, image and radar features are independently extracted, unified into a BEV representation, and subsequently fused by concatenation or cross-attention, as shown in Fig. 1 (a). However, the inherent disparity between the two modalities poses a significant challenge when relying solely on BEV features for fusion. Due to radars’ hardware constraints, e.g., bandwidth and array design, their limited spatial resolution results in sparse radar BEV features. On the other hand, camera BEV features are generated from dense image features, but have the issue of feature distortion due to the inaccurate depth estimation in view transformation [43]. In contrast, the original perspective image features offer a semantically rich representation free from distortions, which has the potential to aid camera-radar fusion. This underscores the necessity of heterogeneous feature fusion across perspectives of the front view and BEV. Inspired by this finding, this work aims to explore an effective cross-perspective fusion framework that can accommodate resolution and semantic discrepancies between the two modalities.
Motivated by the above analysis, we ask one essential question: What kind of fusion paradigm can remain unaffected by feature density while effectively utilizing information from different views? Back to the detection algorithms, the query-based approach [4] shows potential in addressing this dilemma. Specifically, the object query initialized in the 3D space can be leveraged as a medium for abstracting features from arbitrary projection views.
@@
Formally, we propose RaCFormer, a query-based radar-camera fusion framework that improves camera-radar fusion by sampling object-relevant features from both perspective and bird’s-eye views, as illustrated in the dual-view fusion paradigm in Fig. 1 (b). Our framework has three main designs: linearly increasing circular query initialization, radar-aware depth prediction, and temporal radar BEV encoding. The first optimizes the initialization distribution of queries, while the latter two refine and bolster the BEV representation. Specifically, we propose a circular query initialization strategy that places query points along concentric circles to align the projection principle of sensors. Additionally, we ensure a linear increase in the number of queries from inner to outer circles, thereby mitigating the issue of queries being much sparser at distant ranges compared to nearby areas. Furthermore, conventional automotive radars exhibit significant height estimation errors due to their limited vertical angular resolution. Therefore, we assign a default height and project radar points onto the image features to enhance depth prediction for view transformation, refining the camera BEV features. We also utilize the radar’s Doppler effect to track object velocities by employing an implicit dynamic catcher with convolutional gated recurrent units, effectively capturing temporal elements on multi-frame radar BEV features.
@@
  • We perform experiments on the nuScenes and VoD datasets. Our method achieves state-of-the-art performance on both the nuScenes test set and the VoD dataset.
@@ 3.1 Overall Framework
RaCFormer, as shown in Fig. 2, offers a query-based 3D object detection framework that integrates radar and camera inputs. The core modules of the framework include an image encoder, a pillar encoder, a radar-aware depth head, an LSS view transformation module, an implicit dynamic catcher, and a transformer decoder. The image encoder extracts features from camera frames, while the pillar encoder processes radar points and flattens the features to BEV. Subsequently, the radar BEV features are refined by an implicit dynamic catcher to capture moving elements. Radar points are also projected into the image plane and combined with visual features in the radar-aware depth head to form a depth probability distribution $D^{\prime}$. The enhanced depth distribution is merged with image features in the LSS module to generate camera BEV features. Queries serve as a medium for cross-perspective and cross-modality feature fusion, initialized in an adjustable circular distribution and refined by the transformer decoder. The transformer decoder, comprising $L$ layers, includes a scale-adaptive self-attention module [31] for dynamically adjusting receptive fields, two ray-sampling modules for extracting BEV and image-view features, and an adaptive mixer [9] for feature aggregation. Finally, the classification and regression heads interpret the refined queries for accurate object detection.
@@ 3.2 Camera-transformed BEV Feature Generation with Radar-aware Depth Prediction
Radar-aware Depth Prediction: We propose to enhance the image features using radar data for better depth estimation. However, conventional automotive radars provide distance and velocity measurements within their field of view, but their limited vertical angular resolution leads to significant height estimation errors. As depicted in Fig. 3(a), many radar points are projected onto the image with their vertical coordinates falling outside the objects’ 2D bounding boxes, due to inaccuracies in the z-coordinates of the raw radar points. Therefore, we design a pre-processing step before the depth head, shown in Fig. 3(b). To maximize the number of radar points falling onto the image’s field of view, we initially set $z_{r}=1$ for all points, denoted as $(x_{r},y_{r},z_{r})$, and then project them onto the image plane using the camera’s intrinsic parameters. The specific transformation formula is as follows:
@@ 4.1 Datasets and Metrics
Table 2: Comparison on the nuScenes test set. The VoVNet-99 (V2-99) [21] is pre-trained from DD3D [42] with extra data. ‘L’, ‘C’, and ‘R’ represent LiDAR, camera, and radar, respectively. “(+$t$)" indicates using future and historical frames, each by $t$ frames.
4.2 Implementation Details
@@ 4.4 Ablation Studies
Feature Decoding and Fusion: Tab. 4 illustrates the impact of feature decoding methods and perspective selection. The first two rows present the results of employing the standard BEVFusion paradigm [34] using the concatenation operation or deformable cross-attention for BEV feature fusion, combined with a center-head decoder. Switching to a transformer decoder with queries, focusing solely on the sampling of BEV features, yields a 1.9% enhancement in mAP and a 2.5% improvement in NDS, along with a significant reduction in mAVE but an increase in mATE. Further sampling of image-view features improves 3.1% at mAP and 4.1% at NDS, respectively.
Radar-aware Depth Prediction: In Tab. 5, we examine the impact of embedding radar points’ depth and RCS value. The model that only incorporates radar depth embedding achieves an improvement of 0.7% in mAP and 1.2% in NDS. Similarly, the embedding of RCS enhances 0.5% mAP and 0.8% NDS. When both embeddings are used, the performance is further improved, with overall gains of 1.1% in mAP and 2.1% in NDS.
@@
Concat. B 35.9 42.7 0.625 0.289 0.613
Def. Attn. B 38.6 45.1 0.576 0.280 0.601
Queries B 40.5 47.6 0.655 0.281 0.365
Queries B+I 43.6 51.7 0.577 0.274 0.341
Table 4: Ablation study of the feature decoding. ‘B’ and ‘I’ correspond to the BEV and image view, respectively.
Implicit Dynamic Catching: In Tab. 6, we evaluate the performance of our implicit dynamic catcher (IDC) by evaluating mAP and mAVE for moving objects in the nuScenes validation set, divided by objects’ velocity: static (0 m/s), slow ($<$5 m/s), and fast ($>$5 m/s) objects. Adding radar data without IDC substantially increases the relative mAP of slow and fast objects by 49.2% and 48.1%, while decreasing relative mAVE by 30.2% and 11.8%. With IDC, RaCFormer further improves relative mAP by 4.3% and 3.5% and relatively reduces mAVE by 2.9% and 2.0%. The integration of radar data and the IDC module reveals subtle temporal feature changes, leading to more pronounced optimization on slow-moving objects.
@@
43.6 51.7 0.577 0.274 0.341
44.3 52.9 0.560 0.268 0.319
44.1 52.5 0.564 0.269 0.332
44.7 53.8 0.558 0.270 0.313
Table 5: Ablation study about the depth and RCS embedding in the pre-processing step of radar-aware depth prediction.
Input IDC 0 m/s (0, 5] m/s $>$5 m/s
@@
Table 9: Analysis of robustness using Car class AP. “All" denotes that the single modality is entirely off.
Inference Time: To enable real-time detection, we develop a lightweight version of our model: it uses 4 historical frames, 450 queries, and a BEV grid of $64\times 64$ with a resolution of 1.6 meters per voxel. These adjustments reduce the computational load compared to our standard model. Despite these reductions, our model still delivers state-of-the-art mAP and NDS of 51.0% and 58.8%, respectively, outperforming HyDRa [48] by 1.6% in mAP and 2.8% in NDS. Operating on a single RTX 3090 GPU, it achieves a frame rate of 12 FPS, satisfying real-time requirements.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62332016) and the Key Research Program of Frontier Sciences, CAS (No. ZDBS-LY-JSC001).